consider a call option and put option on the same underlying stock with the same exercise price and time to maturity.the call price is 2.59,the underlying stock price is 28.63,the exercise price on both options is 26.18,the risk-free rate is6.21%,the time to maturity on both options is 0.47 years and the stock pays a 1.64ishare divident in 0.28 years ,determine the price of the put price now.
Answer #1

The current question is based on Put Call parity concept,

P0+ S0   = C0 + (D*e-r*t + X*e-r*t)

=> P0   = C0 + (D*e-r*t + X*e-r*t) - S0

here, given data

P0 = 2.59 + (1.64 * e-0.0621*0.28 + 26.18 * e-0.0621*0.47 ) - 28.63

P0 = 0.99535 = $ 1.00